OpenStudy (anonymous):

help! medals rewarded What are the amplitude, period, and phase shift of the given function? f(t)=-1/2sin(4t-2pi)

OpenStudy (freckles):

the amplitude should be the easiest number to spot in these

OpenStudy (freckles):

f(x)=a*sin(bt+c) The number a tells us the height the function will stretch from the graph's "middle axis" (this is the axis I like to call middle because it is in the center of the graph) For f(x)=sin(x) the "middle axis" is y=0 and from it the graph stretches up 1 and stretches down 1 this is because the number in front of sin(x) is 1 (it would have the same amplitude if it was f(x)=-sin(x) What I'm saying it the amplitude is simply |a| Now the period? the period of y=sin(x) is 2pi so the period of y=sin(bx) is 2pi/b Phase shift is where the graph starts one of its first periods? This number can be found by solving the inside=0. That is in your case 4t-2pi=0

OpenStudy (freckles):

Summary: if y=asin(bt+c) amp=|a| period=2pi/b phase shift comes from solve bt+c=0

OpenStudy (freckles):

what do you think the amp is?

OpenStudy (anonymous):

1/2

OpenStudy (freckles):

ok cool period?

OpenStudy (anonymous):

is it 1/2pi

OpenStudy (freckles):

2pi/4=1pi/2=pi/2 yeah

OpenStudy (freckles):

phase shift?

OpenStudy (anonymous):

the same

OpenStudy (freckles):

lol yeah also the reason the phase shift can be found from setting the inside=0 is because y=sin(x) has phase shift at x=0 that is its inside=0

OpenStudy (freckles):

you can use the same information I gave you above for y=acos(bt+c)

OpenStudy (anonymous):

thanks so much

OpenStudy (freckles):

np